# Analog Computing for Signal Processing and Communications – Part II: Toward Gigantic MIMO Beamforming

### v2

@@ Abstract
Index Terms:
Analog beamforming, analog computing, massive multiple-input multiple-output (MIMO), minimum mean square error (MMSE), zero-forcing (ZF).
@@ I Introduction
In Part I of this paper, we have reviewed the growing interest in analog computing, driven by advancements in several technologies such as optical systems, metamaterials, and resistive memory arrays. Digital computing, despite its dominance, faces fundamental limitations in speed and power consumption, mainly due to the reliance on analog-to-digital converters (ADCs) and digital-to-analog converters (DACs). Analog computing addresses these challenges by enabling energy-efficient and highly parallelized computations at the speed of light. While Part I has focused on how analog computing can be leveraged for signal processing operations, Part II explores how these principles can be applied to achieve efficient and flexible beamforming in the analog domain, paving the way for future wireless communications.
@@
To address the high hardware cost associated with digital beamforming, alternative strategies such as analog beamforming and hybrid analog-digital beamforming have been proposed. Analog beamforming steers the beam pattern by solely requiring a single RF chain [5]. This is achieved by connecting the RF chain to the antennas via phase shifters or time delay elements, which impose a constant modulus constraint on the beamforming weights. Thus, while analog beamforming offers low hardware complexity, it is characterized by limited flexibility. To balance hardware cost and flexibility by bridging analog and digital beamforming, hybrid analog-digital beamforming techniques have been widely investigated [5, 6, 7, 8, 9]. In hybrid beamforming, the beamforming process is divided into two stages. First, the transmitted symbols are digitally processed in the baseband domain with a low-dimensional digital beamforming matrix, and fed to a reduced number of RF chains. Second, the RF chains are connected to the transmitting antennas through a network of phase shifters, operating analog beamforming in the RF domain. The reversed process is applied at the receiver.
Another method to efficiently perform beamforming in the analog domain is given by reconfigurable intelligent surface (RIS) [10, 11]. This technology has recently emerged enabling the control of the wireless channel through surfaces made of multiple passive elements with reconfigurable scattering properties. Beamforming in the analog domain can be achieved through RIS by deploying a RIS in close proximity to an active transceiving device equipped with a reduced number of antennas. In this way, the effective radiation pattern of the active device can be optimized by reconfiguring the RIS elements [12, 13, 14]. Two extensions of this technology have been proposed to achieve additional flexibility over conventional RIS. First, beyond diagonal RIS (BD-RIS) include more general RIS architectures, characterized by a scattering matrix not restricted to be diagonal [15, 16, 17]. By deploying a BD-RIS close to an active device, its additional flexibility over conventional RIS in manipulating the radiation pattern can enable even higher performance [18]. Second, stacked intelligent metasurface (SIM) technology considers multiple stacked transmissive RISs deployed close to the active device [19, 20, 21]. The additional flexibility provided by the multiple layers has shown increased performance over just deploying a single conventional RIS.
@@
In addition to alleviating the hardware cost, analog-domain strategies have been also explored to reduce the computational complexity of future wireless networks, by performing specific mathematical operations in the analog domain. To this end, in-memory computing has been proposed to accelerate ridge regression with applications to massive MIMO [29]. SIM has been used to compute the discrete Fourier transform (DFT) of the incident signal in the analog domain, which is particularly beneficial for direction-of-arrival (DoA) estimation [30]. Moreover, over-the-air computing emerged as a form of analog computing that leverages the superposition principle characterizing the electromagnetic (EM) propagation over the wireless channels. Over-the-air computing allows efficient computation of the so-called nomographic functions [31, 32, 33], with impactful applications on federated learning [34].
@@
Numerous technologies have been proposed to reduce the hardware cost of future wireless transceivers through analog beamforming [5]-[28]. However, operating in the analog domain typically decreases the flexibility compared to digital beamforming, with a consequent reduction in performance. Thus, whether it is possible to maintain maximum performance while simultaneously minimizing hardware cost and computational complexity through analog computing [29]-[34] remains an open question. In Part II of this paper, we address this gap by introducing a novel beamforming strategy enabled by microwave linear analog computers (MiLACs), denoted as MiLAC-aided beamforming. This beamforming strategy maintains the same flexibility as digital beamforming while operating entirely in the analog domain. Because of its fully analog nature, it minimizes the number of RF chains and relies only on low-resolution ADCs/DACs, reducing the hardware cost. By leveraging the computational capabilities of MiLACs, it also minimizes the computational complexity required for beamforming. More in detail, the contributions of Part II of this paper are as follows.
@@
First, we demonstrate that a MiLAC can efficiently compute five mathematical operations of practical interest in MIMO communications, such as least squares (LS) and matched filtering (MF). To this end, we first show how these five operations can be seen as special cases of the linear minimum mean square error (LMMSE) estimator for linear observation processes, and then discuss how the microwave network of the MiLAC needs to be reconfigured to compute them. These operations can be computed with a MiLAC with significantly reduced computational complexity compared to conventional digital computing.
@@
Second, we propose MiLAC-aided beamforming as a new beamforming strategy enabled by MiLAC scalable to gigantic MIMO systems. MiLAC-aided beamforming can perform precoding at the transmitter and combining at the receiver with arbitrary matrices, i.e., with the same flexibility as digital beamforming. While having the maximum flexibility, MiLAC-aided beamforming allows us to perform specific operations in the analog domain, such as zero-forcing (ZF) precoding at the transmitter and ZF combining at the receiver, with important gains in terms of computational complexity. In addition, it allows to precode or combine a signal with a fixed matrix, such as the DFT matrix, entirely in the analog domain, requiring no digital operations.
@@
Notation: Please, refer to the notation introduced in Section I of Part I of this paper.
Figure 1: Representation of a $P$-port MiLAC with input on $N$ ports, where $P=N+M$.
@@ II Microwave Linear Analog Computer
In this section, we briefly review the concept of MiLAC, defined as an analog computer that linearly processes microwave signals, as introduced in Part I of this paper. In general, a MiLAC can be seen as a $P$-port reconfigurable microwave network, which can be characterized by its admittance matrix $\mathbf{Y}\in\mathbb{C}^{P\times P}$. We assume this microwave network to be made of $P^{2}$ tunable admittance components interconnecting each port to ground and to all the other ports, such that the microwave network can be reconfigured to assume any arbitrary admittance matrix $\mathbf{Y}$. The admittance component interconnecting port $k$ to ground is denoted as $Y_{k,k}\in\mathbb{C}$, for $k=1,\ldots,P$, while the admittance component interconnecting port $k$ to port $i$ is denoted as $Y_{i,k}\in\mathbb{C}$, for $\forall i\neq k$. The input signal $\mathbf{u}=[u_{1},\ldots,u_{N}]^{T}\in\mathbb{C}^{N\times 1}$ is applied by $N$ voltage generators with series admittance $Y_{0}$ to the first $N$ ports of the MiLAC, with $N\leq P$, as represented in Fig. 1. The output signal is read on all $P$ ports, where $P=N+M$ with $M$ denoting the number of ports with no input, and is denoted by $\mathbf{v}=[v_{1},\ldots,v_{P}]^{T}\in\mathbb{C}^{P\times 1}$, as shown in Fig. 1. For convenience of representation, we introduce the output vectors $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ as $\mathbf{v}_{1}=[v_{1},\ldots,v_{N}]^{T}\in\mathbb{C}^{N\times 1}$ and $\mathbf{v}_{2}=[v_{N+1},\ldots,v_{p}]^{T}\in\mathbb{C}^{M\times 1}$, such that $\mathbf{v}=[\mathbf{v}_{1}^{T},\mathbf{v}_{2}^{T}]^{T}$. We consider a MiLAC with $M>0$ ports with no input in the following.
It has been shown in Part I of this paper that the output vectors $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ can be expressed as a function of the admittance matrix $\mathbf{Y}$ and the input $\mathbf{u}$ by introducing a matrix $\mathbf{P}\in\mathbb{C}^{P\times P}$ as
@@
for $i,k=1,\ldots,P$, as discussed in Part I of this paper. Remarkably, a MiLAC can efficiently compute (3)-(4) and (5)-(6) in the analog domain, while they would require computationally expensive matrix-matrix products and matrix inversion operations if digitally computed.
@@ III-A LMMSE Estimator for Linear Observation Processes
Consider a linear observation process $\mathbf{y}=\mathbf{H}\mathbf{x}+\mathbf{n}$, where $\mathbf{y}\in\mathbb{C}^{Y\times 1}$ is the observation vector, $\mathbf{H}\in\mathbb{C}^{Y\times X}$ is a known constant matrix, $\mathbf{x}\in\mathbb{C}^{X\times 1}$ is the unknown random vector with mean $\bar{\mathbf{x}}=\mathbf{0}_{X\times 1}$ and covariance matrix $\mathbf{C}_{\mathbf{x}}$, and $\mathbf{n}\in\mathbb{C}^{Y\times 1}$ is the random noise vector with mean $\bar{\mathbf{n}}=\mathbf{0}_{Y\times 1}$ and covariance matrix $\mathbf{C}_{\mathbf{n}}$. We assume that $\mathbf{C}_{\mathbf{x}}$ and $\mathbf{C}_{\mathbf{n}}$ are known and invertible, and that the cross-covariance matrix of $\mathbf{x}$ and $\mathbf{n}$ is $\mathbf{C}_{\mathbf{x}\mathbf{n}}=\mathbf{0}_{X\times Y}$. As discussed in the part I of this paper, the LMMSE estimator of $\mathbf{x}$ is equivalently given by
@@ III-B3 Regularized Least Squares
respectively. Interestingly, the expressions in (13) and (14) are widely used in MIMO wireless communications to realize the so-called regularized zero-forcing beamforming (R-ZFBF) and the minimum mean square error (MMSE) receiver, as effective linear transmission and reception techniques, respectively [35, Chapters 7, 12].
@@ III-B4 Ordinary Least Squares
respectively. Interestingly, OLS is a special case of both GLS and RLS as it includes both their assumptions on $\mathbf{C}_{\mathbf{x}}$ and $\mathbf{C}_{\mathbf{n}}$. As discussed for the GLS, since only one matrix among $\mathbf{H}^{H}\mathbf{H}$ and $\mathbf{H}\mathbf{H}^{H}$ may be invertible depending on the dimensions of $\mathbf{H}$, only one among $\hat{\mathbf{x}}_{\text{OLS},1}$ and $\hat{\mathbf{x}}_{\text{OLS},2}$ may exist, unless $\mathbf{H}$ is a square matrix. Remarkably, the expressions of the OLS are widely employed in MIMO wireless communications to implement zero-forcing beamforming (ZFBF) and the ZF receiver [35, Chapters 7, 12].
@@ III-B5 Ordinary Matched Filtering
Remarkably, OMF is a special case of both GMF and RLS as it jointly accounts for their assumptions on $\mathbf{C}_{\mathbf{x}}$ and $\mathbf{C}_{\mathbf{n}}$. In addition, it is utilized in MIMO wireless communications to realize matched beamforming (MBF) and the MF receiver [35, Chapters 7, 12].
@@
LMMSE $\left(\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{H}+\mathbf{C}_{\mathbf{x}}^{-1}\right)^{-1}\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{u}$ [±$\begin{bmatrix}\pm\mathbf{C}_{\mathbf{n}}&\mathbf{H}\\ \mathbf{H}^{H}&\mp\mathbf{C}_{\mathbf{x}}^{-1}\end{bmatrix}$
GLS $\left(\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{H}\right)^{-1}\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{u}$ [±$\begin{bmatrix}\pm\mathbf{C}_{\mathbf{n}}&\mathbf{H}\\ \mathbf{H}^{H}&\mathbf{0}_{X}\end{bmatrix}$
GMF $\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{u}$ [±$\begin{bmatrix}\pm\mathbf{C}_{\mathbf{n}}&\mathbf{0}_{Y\times X}\\ \mathbf{H}^{H}&\mp\mathbf{C}_{\mathbf{x}}^{-1}\end{bmatrix}$
RLS $\left(\mathbf{H}^{H}\mathbf{H}+\lambda\mathbf{I}_{X}\right)^{-1}\mathbf{H}^{H}\mathbf{u}$ [±$\begin{bmatrix}\pm\mathbf{I}_{Y}&\mathbf{H}\\ \mathbf{H}^{H}&\mp\lambda\mathbf{I}_{X}\end{bmatrix}$
OLS $\left(\mathbf{H}^{H}\mathbf{H}\right)^{-1}\mathbf{H}^{H}\mathbf{u}$ [±$\begin{bmatrix}\pm\mathbf{I}_{Y}&\mathbf{H}\\ \mathbf{H}^{H}&\mathbf{0}_{X}\end{bmatrix}$
OMF $\lambda^{-1}\mathbf{H}^{H}\mathbf{u}$ [±$\begin{bmatrix}\pm\mathbf{I}_{Y}&\mathbf{0}_{Y\times X}\\ \mathbf{H}^{H}&\mp\lambda\mathbf{I}_{X}\end{bmatrix}$

### v3

@@ I Introduction
In Part I of this paper [2], we have reviewed the growing interest in analog computing, driven by advancements in several technologies such as optical systems, metamaterials, and resistive memory arrays. Digital computing, despite its dominance, faces fundamental limitations in speed and power consumption, mainly due to the reliance on analog-to-digital converters (ADCs) and digital-to-analog converters (DACs). Analog computing addresses these challenges by enabling energy-efficient and highly parallelized computations at the speed of light. While Part I has focused on how analog computing can be leveraged for signal processing operations, Part II explores how these principles can be applied to achieve efficient and flexible beamforming in the analog domain, paving the way for future wireless communications.
@@
To address the high hardware cost associated with digital beamforming, alternative strategies such as analog beamforming and hybrid analog-digital beamforming have been proposed. Analog beamforming steers the beam pattern by solely requiring a single RF chain [6]. This is achieved by connecting the RF chain to the antennas via phase shifters or time delay elements, which impose a constant modulus constraint on the beamforming weights. Thus, while analog beamforming offers low hardware complexity, it is characterized by limited flexibility. To balance hardware cost and flexibility by bridging analog and digital beamforming, hybrid analog-digital beamforming techniques have been widely investigated [6, 7, 8, 9, 10, 11]. In hybrid beamforming, the beamforming process is divided into two stages. First, the transmitted symbols are digitally processed in the baseband domain with a low-dimensional digital beamforming matrix, and fed to a reduced number of RF chains. Second, the RF chains are connected to the transmitting antennas through a network of phase shifters, operating analog beamforming in the RF domain. The reversed process is applied at the receiver.
Another method to efficiently perform beamforming in the analog domain is given by reconfigurable intelligent surface (RIS) [12, 13]. This technology has recently emerged enabling the control of the wireless channel through surfaces made of multiple passive elements with reconfigurable scattering properties. Beamforming in the analog domain can be achieved through RIS by deploying a RIS in close proximity to an active transceiving device equipped with a reduced number of antennas. In this way, the effective radiation pattern of the active device can be optimized by reconfiguring the RIS elements [14, 15, 16]. Two extensions of this technology have been proposed to achieve additional flexibility over conventional RIS. First, beyond-diagonal RIS (BD-RIS) include more general RIS architectures, characterized by a scattering matrix not restricted to be diagonal [17, 18, 19]. By deploying a BD-RIS close to an active device, its additional flexibility over conventional RIS in manipulating the radiation pattern can enable even higher performance [20]. Second, stacked intelligent metasurface (SIM) technology considers multiple stacked transmissive RISs deployed close to the active device [21, 22, 23]. The additional flexibility provided by the multiple layers has shown increased performance over just deploying a single conventional RIS. Reconfigurable metasurfaces similar to those in RIS and SIM have been proposed to enable dynamic metasurface antennas (DMAs), and perform beamforming and combining in the analog domain [24, 25, 26].
@@
In addition to alleviating the hardware cost, analog-domain strategies have been also explored to reduce the computational complexity of future wireless networks by performing specific mathematical operations in the analog domain. To this end, in-memory computing has been proposed to accelerate ridge regression with applications to massive MIMO [34, 35, 36]. While in-memory computing approaches can efficiently process baseband signals, other strategies have emerged to directly perform analog processing at RF. RIS and SIM have been used to compute the discrete Fourier transform (DFT) of the incident signal in the analog domain, which is particularly beneficial for direction-of-arrival (DoA) estimation [37, 38, 39]. Moreover, over-the-air computing emerged as a form of analog computing that leverages the superposition principle characterizing the electromagnetic (EM) propagation over the wireless channels. Over-the-air computing allows efficient computation of the so-called nomographic functions [40, 41, 42], with impactful applications on federated learning [43].
@@
Numerous technologies have been proposed to reduce the hardware cost of future wireless transceivers through analog beamforming [6]-[33] and analog computing [34]-[43]. However, operating in the analog domain typically decreases the flexibility compared to digital beamforming, with a consequent reduction in performance. Thus, whether it is possible to maintain maximum performance while simultaneously minimizing hardware cost and computational complexity through analog computing remains an open question. In Part II of this paper, we address this gap by introducing a novel beamforming strategy enabled by microwave linear analog computers (MiLACs), denoted as MiLAC-aided beamforming. This beamforming strategy maintains the same flexibility as digital beamforming while operating entirely in the analog domain. Because of its fully analog nature, it minimizes the number of RF chains and relies only on low-resolution ADCs/DACs, reducing the hardware cost. By leveraging the computational capabilities of MiLACs, it also minimizes the computational complexity required for beamforming. More in detail, the contributions of Part II of this paper are as follows.
@@
First, we demonstrate that a MiLAC can efficiently process microwave signals to compute five mathematical operations of practical interest in MIMO communications, such as least squares (LS) and matched filtering (MF). To this end, we first show how these five operations can be seen as special cases of the linear minimum mean square error (LMMSE) estimator for linear observation processes, and then discuss how the microwave network of the MiLAC needs to be reconfigured to compute them. These operations can be computed with a MiLAC with significantly reduced computational complexity compared to conventional digital computing.
@@
Second, we propose MiLAC-aided beamforming as a new beamforming strategy enabled by MiLAC scalable to gigantic MIMO systems. MiLAC-aided beamforming can perform precoding at the transmitter and combining at the receiver, directly operating with microwave signals, and maintaining the same flexibility as digital beamforming. While having the maximum flexibility, MiLAC-aided beamforming allows us to perform specific operations in the analog domain, such as zero-forcing (ZF) precoding at the transmitter and ZF combining at the receiver, with important gains in terms of computational complexity. In addition, it allows to precode or combine a signal with a fixed matrix, such as the DFT matrix, entirely in the analog domain, requiring no digital operations.
@@
Notation: Please, refer to the notation introduced in Section I of Part I of this paper [2].
Figure 1: Representation of a $P$-port MiLAC with input on NN ports, where $P=N+M$.
@@ II Microwave Linear Analog Computer
In this section, we briefly review the concept of MiLAC, defined as an analog computer that linearly processes microwave signals, as introduced in Part I of this paper [2]. A MiLAC can be seen as a $P$-port reconfigurable microwave network, which can be characterized through its impedance, admittance, or scattering matrix according to multiport network theory. Since these three representations are equivalent, we chose the admittance matrix $\mathbf{Y}\in\mathbb{C}^{P\times P}$ since its entries are closely related to the tunable components in the MiLAC, as seen in Part I. We assume this microwave network to be made of $P^{2}$ tunable admittance components interconnecting each port to ground and to all the other ports. The admittance component interconnecting port $k$ to ground is denoted as $Y_{k,k}\in\mathbb{C}$, for $k=1,\ldots,P$, while the admittance component interconnecting port $k$ to port $i$ is denoted as $Y_{i,k}\in\mathbb{C}$, for $\forall i\neq k$. Depending on the admittance values $\{Y_{i,k}\}_{i,k=1}^{P}$, the entries of the admittance matrix $\mathbf{Y}$ are give by
$\left[\mathbf{Y}\right]_{i,k}=\begin{cases}-Y_{i,k}&i\neq k\\ \sum_{p=1}^{P}Y_{p,k}&i=k\end{cases},$ (1)
for $i,k=1,\ldots,P$, as derived in Part I. The relationship in (1) shows that if the $P^{2}$ tunable admittance components are arbitrarily reconfigurable, then the admittance matrix $\mathbf{Y}$ can also be arbitrarily reconfigured.
The input signal $\mathbf{u}=[u_{1},\ldots,u_{N}]^{T}\in\mathbb{C}^{N\times 1}$ is applied by $N$ voltage generators with series admittance $Y_{0}$ to the first $N$ ports of the MiLAC, with $N\leq P$, as represented in Fig. 1. The output signal is read on all $P$ ports, where $P=N+M$ with $M$ denoting the number of ports with no input, and is denoted by $\mathbf{v}=[v_{1},\ldots,v_{P}]^{T}\in\mathbb{C}^{P\times 1}$, as shown in Fig. 1. For convenience of representation, we introduce the output vectors $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ as $\mathbf{v}_{1}=[v_{1},\ldots,v_{N}]^{T}\in\mathbb{C}^{N\times 1}$ and $\mathbf{v}_{2}=[v_{N+1},\ldots,v_{p}]^{T}\in\mathbb{C}^{M\times 1}$, such that $\mathbf{v}=[\mathbf{v}_{1}^{T},\mathbf{v}_{2}^{T}]^{T}$. We consider a MiLAC with $M>0$ ports with no input in the following.
It has been shown in Part I that the output vectors $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ can be expressed as a function of the admittance matrix $\mathbf{Y}$ and the input $\mathbf{u}$ by introducing a matrix $\mathbf{P}\in\mathbb{C}^{P\times P}$ as
@@
for $i,k=1,\ldots,P$, as discussed in Part I. Remarkably, a MiLAC can efficiently compute (4)-(5) and (6)-(7) in the analog domain once the input signals are applied at the input ports, while they would require computationally expensive matrix-matrix products and matrix inversion operations if digitally computed.
@@ III-A LMMSE Estimator for Linear Observation Processes
Consider a linear observation process $\mathbf{y}=\mathbf{H}\mathbf{x}+\mathbf{n}$, where $\mathbf{y}\in\mathbb{C}^{Y\times 1}$ is the observation vector, $\mathbf{H}\in\mathbb{C}^{Y\times X}$ is a known constant matrix, $\mathbf{x}\in\mathbb{C}^{X\times 1}$ is the unknown random vector with mean $\bar{\mathbf{x}}=\mathbf{0}_{X\times 1}$ and covariance matrix $\mathbf{C}_{\mathbf{x}}$, and $\mathbf{n}\in\mathbb{C}^{Y\times 1}$ is the random noise vector with mean $\bar{\mathbf{n}}=\mathbf{0}_{Y\times 1}$ and covariance matrix $\mathbf{C}_{\mathbf{n}}$. We assume that $\mathbf{C}_{\mathbf{x}}$ and $\mathbf{C}_{\mathbf{n}}$ are known and invertible, and that the cross-covariance matrix of $\mathbf{x}$ and $\mathbf{n}$ is $\mathbf{C}_{\mathbf{x}\mathbf{n}}=\mathbf{0}_{X\times Y}$. As discussed in part I, the LMMSE estimator of $\mathbf{x}$ is equivalently given by
@@ III-B3 Regularized Least Squares
respectively. Interestingly, the expressions in (14) and (15) are widely used in MIMO wireless communications to realize the so-called regularized zero-forcing beamforming (R-ZFBF) and the minimum mean square error (MMSE) receiver, as effective linear transmission and reception techniques, respectively [44, Chapters 7, 12].
@@ III-B4 Ordinary Least Squares
respectively. Interestingly, OLS is a special case of both GLS and RLS as it includes both their assumptions on $\mathbf{C}_{\mathbf{x}}$ and $\mathbf{C}_{\mathbf{n}}$. As discussed for the GLS, since only one matrix among $\mathbf{H}^{H}\mathbf{H}$ and $\mathbf{H}\mathbf{H}^{H}$ may be invertible depending on the dimensions of $\mathbf{H}$, only one among $\hat{\mathbf{x}}_{\text{OLS},1}$ and $\hat{\mathbf{x}}_{\text{OLS},2}$ may exist, unless $\mathbf{H}$ is a square matrix. Remarkably, the expressions of the OLS are widely employed in MIMO wireless communications to implement zero-forcing beamforming (ZFBF) and the ZF receiver [44, Chapters 7, 12].
@@ III-B5 Ordinary Matched Filtering
Remarkably, OMF is a special case of both GMF and RLS as it jointly accounts for their assumptions on $\mathbf{C}_{\mathbf{x}}$ and $\mathbf{C}_{\mathbf{n}}$. In addition, it is utilized in MIMO wireless communications to realize matched beamforming (MBF) and the MF receiver [44, Chapters 7, 12].
@@
LMMSE $\left(\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{H}+\mathbf{C}_{\mathbf{x}}^{-1}\right)^{-1}\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{u}$ $\begin{bmatrix}\pm\mathbf{C}_{\mathbf{n}}&\mathbf{H}\\ \mathbf{H}^{H}&\mp\mathbf{C}_{\mathbf{x}}^{-1}\end{bmatrix}$
GLS $\left(\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{H}\right)^{-1}\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{u}$ $\begin{bmatrix}\pm\mathbf{C}_{\mathbf{n}}&\mathbf{H}\\ \mathbf{H}^{H}&\mathbf{0}_{X}\end{bmatrix}$
GMF $\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{u}$ $\begin{bmatrix}\pm\mathbf{C}_{\mathbf{n}}&\mathbf{0}_{Y\times X}\\ \mathbf{H}^{H}&\mp\mathbf{C}_{\mathbf{x}}^{-1}\end{bmatrix}$
RLS $\left(\mathbf{H}^{H}\mathbf{H}+\lambda\mathbf{I}_{X}\right)^{-1}\mathbf{H}^{H}\mathbf{u}$ $\begin{bmatrix}\pm\mathbf{I}_{Y}&\mathbf{H}\\ \mathbf{H}^{H}&\mp\lambda\mathbf{I}_{X}\end{bmatrix}$
OLS $\left(\mathbf{H}^{H}\mathbf{H}\right)^{-1}\mathbf{H}^{H}\mathbf{u}$ $\begin{bmatrix}\pm\mathbf{I}_{Y}&\mathbf{H}\\ \mathbf{H}^{H}&\mathbf{0}_{X}\end{bmatrix}$
OMF $\lambda^{-1}\mathbf{H}^{H}\mathbf{u}$ $\begin{bmatrix}\pm\mathbf{I}_{Y}&\mathbf{0}_{Y\times X}\\ \mathbf{H}^{H}&\mp\lambda\mathbf{I}_{X}\end{bmatrix}$